OpenStudy (anonymous):

Find the equation of the line normal to the curve y=secx+x at the point (0,1)

OpenStudy (anonymous):

i dont understand what should i do first ??

OpenStudy (anonymous):

the normal slope is opposite reciprocal of the tangent slope. So just find the tangent slope and then use point slope form

OpenStudy (anonymous):

what should i do ?

OpenStudy (anonymous):

do you know how to take derivative?

OpenStudy (anonymous):

\[\frac{ dy }{ dx }= \sec x \tan x + 1\] put x=0, we get slope of tangent equal to '1'. so the slope of normal is equal to (-1/1)=-1 hence equation of normal is \[(y-1)=-1(x) \] so the equation of normal is \[y=1-x\]

OpenStudy (jdoe0001):

if I read that correctly, you're just being asked to find the equation of the line tangent to sec(x)+x which means the equation of the tangent line to it, which means as sourwing said, find the slope, then use the given point of (0,1) to get the equation of the line

OpenStudy (anonymous):

ok so the slope is and i dont understand what u did after that

OpenStudy (anonymous):

find the derivative. That's what you need to do. Do you know how?

OpenStudy (anonymous):

yep is it secxtanx+1?

OpenStudy (anonymous):

whats next?

OpenStudy (anonymous):

plug 0 in for x

OpenStudy (jdoe0001):

\(\bf \cfrac{d}{dx}[sec(x)+x]\implies f'(x)=sec(x)tan(x)+1\qquad ({\color{blue}{ 0}},1) \\ \quad \\ f'(x)=sec({\color{blue}{ 0}})tan({\color{blue}{ 0}})+1\)

OpenStudy (jdoe0001):

that'd give you the value of the slope recall that a derivative is nothing but the EQUATION to find the slope

OpenStudy (anonymous):

which equals 1. and so the normal slope is -1/1 = -1 answer: y - 0 = -1(x-1)

OpenStudy (jdoe0001):

\(\bf f'(x)=sec({\color{blue}{ 0}})tan({\color{blue}{ 0}})+1\implies f'(x)=\cfrac{1}{cos(0)}\cdot \cfrac{sin(0)}{cos(0)}+1\)

OpenStudy (jdoe0001):

notice your Unit Circle, the values for cosine and sine at 0

OpenStudy (anonymous):

what did u use to show -1/1??

OpenStudy (anonymous):

it's the opposite reciprocal of 1

OpenStudy (jdoe0001):

hhyudg he's using the PERPENDICULAR LINE to THAT TANGENT line, also called as the NORMAL LINE the slope of a line perpendiicular to another is, NEGATIVE RECIPROCAL to that one meaning say if you had a slope of \(\bf \cfrac{1}{{\color{blue}{ 1}}}\qquad reciprocal\implies \cfrac{{\color{blue}{ 1}}}{1}\qquad negative\implies -\cfrac{{\color{blue}{ 1}}}{1}\implies -1\)

OpenStudy (jdoe0001):